OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 -7 7 9 -9

OpenStudy (anonymous):

@kiszer4148

OpenStudy (anonymous):

Do you know how to set this up get rid of one of the variables?

OpenStudy (anonymous):

hmm.. do something on both sides?

OpenStudy (anonymous):

add 3y on both sides?

OpenStudy (anonymous):

you can set it up a few different ways. I will go over the way that i most commonly used. we need to get rid of the variable x from the problem. So we need to multiply the whole top equation by 4 and the whole bottom equation by 5. So you will be left with this 20x+16y=4 and also 20x+15y=-5

OpenStudy (anonymous):

Now you can simply subract the two equations to get: 0x+1y=9 remember that 4-(-5) is 9 so the answer to this problem is a positive 9

OpenStudy (anonymous):

wow! thanks for the help! really grateful! I highly appreciate it! thanks so much! so nice! thanks!